# Quantum dissipative dynamics of driven Duffing oscillator near attractors

### v1

@@ Abstract
We investigate the quantum dissipative dynamics near the stable states (attractors) of a driven Duffing oscillator. A refined perturbation theory that can treat two perturbative parameters with different orders is developed to calculate the quantum properties of Duffing oscillator near the attractors. We obtain the perturbative analytical results for the renormalized level spacings and the quantum-fluctuation-induced distribution of higher energy levels near the attractor, which are then verified by numerical simulations. Furthermore, we demonstrate that strong damping induces a slight renormalization of level spacing, and the renormalized Bose distribution affected by dephasing. Our work provides new insights into the quantum dynamics of driven Duffing oscillator and offers a theoretical framework that can be applied to related quantum system near their stable states.
@@ I Introduction
The driven Duffing oscillator Duf18 , a paradigmatic model for various nonlinear mechanical systems and nonlinear optical phenomena, has been fascinating physicists for a long time by its rich dynamical behaviors such as bistability, bifurcation, and chaotic trajectories. In recent years, the Duffing oscillator has received renewed attention as the quantum regime of nanomechanical oscillators becomes experimentally accessible. The interplay between quantum effects, nonequilibrium dynamics, and nonlinear effects makes the driven Duffing oscillator a model with broad applicability across multiple domains, e.g., mechanical metrology Dyk12m ; Cle02 ; Lif08 ; Poo12 , cavity and circuit quantum electrodynamics Jea24 ; Mav16 ; Boi10 , nano- and opto-mechanics Buk06 ; Alm07 ; Del07 ; Kip11 ; Dyk22 ; Pet14 ; Bol23 ; Hel24 ; Dyk22R , and cold atoms Art03 ; Got19 ; Par12 . Notable examples include bifurcation-based quantum measurement devices, where the low- and high-amplitude states of bistability are entangled with the ground and excited states of qubits respectively, enabling the analysis of the qubit’s state through the detection of classical signals Ald05 ; Sid04 ; Sid06 ; Man07 ; Sid09 .
@@
The dissipative dynamics of driven Duffing oscillators have been extensively studied. In the underdamped regime near a bifurcation point, a scaling law for the noise-induced escape from metastable states was established Dyk05 . It was also revealed that the quantum activation process has distinct temperature dependency compared to that for the quantum tunneling process Dyk07 . The distinct transition rate scaling behaviors near bifurcation points were also revealed in the driven mesoscopic Duffing oscillator Guo10 . It was found that the bifurcation point is shifted by the quantum effect and a linear scaling behavior for the tunneling rate with the driving distance to the shifted bifurcation point Guo11 . Recent advances also showed that the quadrature squeezing can enhance the Wigner negativity in a Duffing oscillator, demonstrating a promising approach to generate nonclassical states in macroscopic mechanical systems Ros24 . For two Duffing oscillators coupled via nonlinear interactions, the stationary paired solutions and their dynamical stability were demonstrated Hel24 . In a coupled system consisting of a time-delayed Duffing oscillator (as a driver system) and a non-delayed Duffing oscillator (as a response system), the phenomenon of transmitted resonance was investigated Coc24 .
@@
In this paper, we focus on investigating quantum dissipative dynamics near the attractors of a driven Duffing oscillator. We develop an effective quantum master equation that can address quantum fluctuations, thermal effects, damping, and dephasing in a unified framework. We quantify the occupation of high levels near the bottom of the potential well by the the quantum squeezing effects. We also demonstrate the effects of strong damping and dephasing on the system’s dynamics, including level spacing renormalization and dephasing-modified Bose distributions. By comparing our theoretical predictions with exact numerical simulations, we demonstrate the accuracy and utility of our proposed framework.
@@ II General Theory
Figure 1: (a) Quasienergy landscape in phase space of a driven Duffing oscillator. The extrema correspond to the high amplitude stable state (M) and the low amplitude stable state (m), while the saddle point (s) denotes an unstable state. (b) Cross-section of the quasienergy potential at $P=0$. Quantum energy levels close to the maximum (M) and minimum (m) of the potential are depicted by blue and red lines, respectively. The unstable saddle point (s) is also indicated.
@@ II.1 Model Hamiltonian
An extensive class of macroscopic physical systems, such as Josephson junctions and nanomechanical oscillators can be modeled by the Duffing oscillator in the presence of a periodic driving force with the system Hamiltonian described by Ser07
@@
Here, parameter $m$ ($\Omega$) describes the mass (frequency) of the oscillator, $\gamma$ gives the nonlinearity of Duffing oscillator, and $F(t)=F_{0}(e^{i\nu t}+e^{-i\nu t})$ describes the periodically driving force with frequency $\nu$. By switching to the rotating frame using the transformation $U(t)=\exp(-i\nu a^{\dagger}at)$, where $a^{\dagger}$ and $a$ are the oscillator raising and lowering operators, and applying the rotating wave approximation (RWA), we obtain a time-independent Hamiltonian
$H=(\delta\omega+\chi)a^{\dagger}a+\chi(a^{\dagger}a)^{2}+\epsilon(a^{\dagger}+a)\,,$ (2)
where $\delta\omega=\hbar(\Omega-\nu)$ is the frequency detuning, $\chi=-3\gamma\hbar^{2}/2(m\Omega)^{2}$ is the scaled dimensionless nonlinearity, and $\epsilon=F_{0}\sqrt{{\hbar}/{2m\Omega}}$ is the scaled driving strength. We then introduce the position operator $Q$ and momentum operator $P$ in the rotating frame via
@@ II.2 Renormalized master equation
The characteristic behavior of the driven Duffing system is the bistability manifesting as two stable states: the low-amplitude state (LAS) and the high-amplitude state (HAS). As depicted in Fig. 1(a), these stable states correspond to the extrema in the quasienergy landscape, which are defined as the attractors in the phase space. The unstable state, known as the saddle point, is located on the separatrix, serving as the boundary dividing the basins of the attractors. When damping is present, the system evolves towards the nearby attractor if its initial state lies within the basin of the attractor. However, due to thermal noise, the system does not remain exactly on the attractor, but forms a probability distribution within its basin of attraction. To describe the dissipative dynamics of our system, we employ a Lindblad form master equation
@@
To study the quantum dynamics near the bottom of each stable state, we first transform the system to the center of the attractor using the displacement operator
@@
with ${\tilde{\alpha}}=[\frac{1}{2}\kappa+i(\delta\omega+\chi+2\chi\alpha|^{2})]\alpha+i\varepsilon.$ By choosing $\alpha$ such that $\tilde{\alpha}=0$, the master equation is simplified into
@@
which has the transformation property $S^{\dagger}aS=va+ua^{\dagger}$ with $v=\cosh(|\xi|)$ and $u=-\frac{|\xi|}{\xi}\sinh(|\xi|)$. Bu defining the squeezed density operator $\bar{\rho}=S^{\dagger}\tilde{\rho}S$, we transform the master equation Eq. (10) into the following form
@@ II.3 Orders of perturbative parameters
To perform perturbation calculations for the renormalized Hamiltonian (II.2), we can choose dimensionless Planck constant $\lambda=-\chi/(4\Delta)$ as the natural choice for the perturbation parameter. However, since the displacement parameter $\alpha$ is also a function of $\lambda$, it is subtle to properly organize the perturbative terms according to their respective orders. In fact, the Hamiltonian (II.2) should be written in different forms for different attractors. The stable state of the driven Duffing oscillator can be approximated as a coherent state $|\alpha\rangle$. By applying the variational principle in quantum mechanics, i.e., $\partial_{\alpha}\langle\alpha|H|\alpha\rangle=0$, we obtain two solutions of the steady coherent number: a smaller one $|\alpha_{l}|^{2}\approx\beta/(2\lambda)$ for the LAS and a larger one $|\alpha_{h}|^{2}\approx 1/(2\lambda)$ for the HAS.
@@
with each term given by
@@
$\frac{\bar{H}^{h}}{\delta\omega}$ $=$ $h^{h}_{0}+\gamma h^{h}_{1}+\gamma h^{h}_{2}$ (26)
with each term given by
@@
The behavior near the bottom of the LAS is relatively simple and can be modeled using a harmonic oscillator. However, for the HAS, the nonlinear term $\chi(a^{\dagger}a)^{2}\approx\chi|\alpha_{h}|^{2}$ becomes prominent, and the oscillator behaves as a highly squeezed coherent state. In the following sections, we will apply our perturbative method to calculate the crucial quantities related to the HAS of the driven Duffing oscillator, namely the level spacing, averaged position of energy level, and the effective temperature in the vicinity of the HAS attractor.
@@ III.1 Quantum dynamics of HAS
The quantum dynamics of the driven Duffing oscillator near the HAS attractor exhibit a rich interplay between nonlinearity, quantum fluctuations, and thermal noise. In this section, we discuss the quantum properties of the HAS using the renormalized master equation combined with a refined perturbation theory.
@@ III.1.1 Level Spacing
Figure 2: Comparison of perturbation calculations and exact results for energy level spacing. The exact results (lowest curve) for level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ are compared with the zeroth-order (constant spacing), the second-order (second highest line, $\gamma^{2}=\lambda$) and the fourth-order (third highest line, $\gamma^{4}=\lambda^{2}$) corrections. Excellent agreement is observed for low level numbers near the potential well bottom, while higher-order corrections are required for large $n$. Parameters: $\lambda=0.016$ and $\beta=4/75$.
@@
The nonlinearity term $\chi(a^{\dagger}a)^{2}$ in the Hamiltonian, having an opposite sign to $\delta\omega$, results in a decrease in the level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ as we approach the saddle point, as illustrated in Fig. 1(b). One can calculate the level spacings with standard perturbation theory by treating the sum of $\gamma h_{1}^{h}$ and $\gamma h_{2}^{h}$ in Eq. (26) as one perturbation term. However, it becomes a challenge to control the accuracy of the level spacings by the perturbative parameter. We find it necessary to distinguish between these two perturbative terms in the perturbation calculations to accurately determine the level spacing. To address this, we have developed a double perturbation theory framework that is particularly suited for the HAS Hamiltonian containing second-order small terms, see the details in Appendix A,
@@
In Fig. 2, we compare our perturbation calculations with the exact numerical results obtained by diagonalizing the original Hamiltonian Eq. (2), which shows an excellent agreement for energy levels near the bottom of the potential well. Under the zeroth-order perturbation approximation, the energy level spacing remains constant across all levels, similar to that of harmonic oscillators. The second-order and fourth-order corrections correspond to accuracies up to $\gamma^{2}=\lambda$ and $\gamma^{4}=\lambda^{2}$, respectively. Higher-order perturbation calculations become necessary for levels farther from the bottom.
@@
Figure 3: Average position of energy levels. Average position $\langle N|a|N\rangle$ of the energy level $N$, illustrating the shifts due to quantum fluctuations. Under the harmonic approximation, $\langle N|a|N\rangle$ is a constant for every level (lowest line). Considering higher-order corrections, we observe changes in $\langle N|a|N\rangle$ with levels. Comparing with the second order perturbation result (highest line), the third order result (second highest line) already align well with the exact one (the bold line).
III.1.2 Averaged level position
@@
We denote the eigenstate of the renormalized Hamiltonian $\bar{H}$ as $|n^{\prime}\rangle$ that is generally a superposition of harmonic oscillator eigenstates $|n^{\prime}\rangle=|n\rangle+\sum_{k\neq n}\xi_{kn}|k\rangle$, where the superposition coefficients $\xi_{kn}$ are provided in the Appendix A. The eigenstate $|N\rangle$ of the original Hamiltonian (2) is related to that of the renormalized Hamiltonian (II.2) via the relationship $|N\rangle=DS|n^{\prime}\rangle$. The matrix element $\langle N|a|M\rangle$ for different levels $|N\rangle$ and $|M\rangle$ is then given by
@@
The matrix element $\langle N|a|N\rangle$ provides insight into the average position of each level in the phase space. Under the harmonic approximation ($\xi_{kn}=0$), $\langle N|a|N\rangle$ is a constant $\alpha$ for all levels. However, considering higher-order corrections, the average position $\langle N|a|N\rangle$ changes with energy level, as depicted in Fig. 3. The perturbation calculations agree well with numerical results.
Figure 4: Comparison of stationary probability distributions. (a) Comparison of the stationary probability distribution obtained through perturbation theory to zeroth order (red line), second order (blue lines) and exact numerical simulations (black line). (b) Relative error $\Delta p_{n}/p_{n}$ in the stationary distribution obtained from perturbation theory compared to numerical results. The discrepancy increases for higher levels but remains less than 1.
@@
Figure 6: Effective temperature of energy levels. The stationary probability distributions $N_{eff}(n)$ of level $|n\rangle$ obtained through perturbation theory to zeroth order (red line) and second order (blue lines) are compared with the exact numerical simulations (black line). The zero-order effective temperature term gives a constant effective temperature. The second order results in changes in $N_{eff}(n)$, which is quite accurate for levels near the bottom.

### v2

@@ Abstract
We investigate the quantum dissipative dynamics near the stable states (attractors) of a driven Duffing oscillator. A refined perturbation theory that can treat two perturbative parameters with different orders is developed to calculate the quantum properties of the Duffing oscillator near the attractors. We obtain the perturbative analytical results, that go beyond the standard linearization approach, for the renormalized level spacings, the orbital displacements, and the effective temperature near the classical attractor. Furthermore, we demonstrate that strong damping induces additional slight renormalization of level spacings, and the Bose distribution together with dephasing. Our work provides new insights into the quantum dynamics of the driven Duffing oscillator and offers a theoretical framework that can be applied to related quantum systems near their stable states.
@@ I Introduction
The driven Duffing oscillator Duf18 , a paradigmatic model for various nonlinear mechanical systems and nonlinear optical phenomena, has fascinated physicists for a long time with its rich dynamical behaviors such as bistability, bifurcation, and chaotic trajectories. In recent years, the Duffing oscillator has received renewed attention as the quantum regime of nanomechanical oscillators becomes experimentally accessible. The interplay between quantum effects, nonequilibrium dynamics, and nonlinear effects makes the driven Duffing oscillator a model with broad applicability across multiple domains, e.g., mechanical metrology Dyk12m ; Cle02 ; Lif08 ; Poo12 , chaotic dynamics Kro23 ; Gho23 , cavity and circuit quantum electrodynamics Zhang13n ; Jea24 ; Mav16 ; Boi10 , nano- and opto-mechanics Buk06 ; Alm07 ; Del07 ; Kip11 ; Dyk22 ; Ken24 ; Pet14 ; Bol23 ; Hel24 ; Dyk22R ; Mil08 , and cold atoms Art03 ; Got19 ; Par12 . Notable examples include bifurcation-based quantum measurement devices, where the low- and high-amplitude states of bistability are entangled with the ground and excited states of qubits, respectively, enabling the analysis of the qubit states through the detection of classical signals Ald05 ; Sid04 ; Sid06 ; Man07 ; Sid09 .
@@
In practice, physical systems inevitably interact with their environment, leading to the decoherence of quantum states and dissipation of energy. The dissipative dynamics of driven Duffing oscillators have been extensively studied Chen23New ; Mar21New . In the underdamped regime near a bifurcation point, a scaling law for the noise-induced escape from metastable states was established Dyk05 . Close to the bottom of the well in a parametrically driven Duffing oscillator, it was identified that the energy dependence of the level spacings captured by the perturbative approach beyond linearization gives rise to a fine structure in the power spectrum Dyk11n . It was also revealed that the quantum activation process has distinct temperature dependency compared to that for the quantum tunneling process Dyk07 . The distinct transition rate scaling behaviors near bifurcation points were also revealed in the driven mesoscopic Duffing oscillator Guo10 . It was found that the bifurcation point is shifted by the quantum effect and a linear scaling behavior for the tunneling rate with the driving distance to the shifted bifurcation point Guo11 . Recent advances also showed that the quadrature squeezing can enhance the Wigner negativity in a Duffing oscillator, demonstrating a promising approach to generate nonclassical states in macroscopic mechanical systems Ros24 . For two Duffing oscillators coupled via nonlinear interactions, the stationary paired solutions and their dynamical stability were demonstrated Hel24 . In a coupled system consisting of a time-delayed Duffing oscillator (as a driver system) and a non-delayed Duffing oscillator (as a response system), the phenomenon of transmitted resonance was investigated Coc24 .
@@
In this paper, we focus on investigating quantum dissipative dynamics near the attractors of a driven Duffing oscillator. We develop an effective quantum master equation that can address quantum fluctuations, thermal effects, damping, and dephasing in a unified framework. Our approach is essential to quantify the occupation of high levels near the bottom of the potential well by the quantum squeezing effects. We also demonstrate the effects of strong damping and dephasing on the system’s dynamics, including level spacing renormalization and dephasing-modified Bose distributions. While prior studies within linearized frameworks have successfully captured phenomena such as effective temperatures Dyk11n ; Pea10n ; Lem15n , the refined perturbation theory presented in our work goes beyond the standard linearization approach allowing us to investigate the energy dependence of the level spacing, the orbital displacement, and the effective temperature. To address this, we develop a refined perturbation theory capable of treating two perturbative parameters with distinct orders, enabling a unified analysis of nonlinear and dissipative quantum effects near the attractors. By comparing our theoretical predictions with exact numerical simulations, we demonstrate the accuracy and utility of our proposed framework.
@@ II.1 Model Hamiltonian
An extensive class of macroscopic physical systems, such as Josephson junctions and nanomechanical oscillators can be modeled by the Duffing oscillator in the presence of a periodic driving force, with the system Hamiltonian described by Ser07
@@
Here, parameter $m$ ($\Omega$) describes the mass (frequency) of the oscillator, $\gamma$ gives the nonlinearity of Duffing oscillator, and $F(t)=F_{0}(e^{i\nu t}+e^{-i\nu t})$ describes the periodical driving force with frequency $\nu$. By switching to the rotating frame using the transformation $U(t)=\exp(-i\nu a^{\dagger}at)$ with $a^{\dagger}$ ($a$) the raising (lowering) operator, and applying the rotating wave approximation (RWA), we obtain a time-independent Hamiltonian
$H=(\delta\omega+\chi)a^{\dagger}a+\chi(a^{\dagger}a)^{2}+\epsilon(a^{\dagger}+a).$ (2)
Here, parameter $\delta\omega=\hbar(\Omega-\nu)$ is the frequency detuning, $\chi=-3\gamma\hbar^{2}/2(m\Omega)^{2}$ is the scaled dimensionless nonlinearity, and $\epsilon=F_{0}\sqrt{{\hbar}/{2m\Omega}}$ is the scaled driving strength. We then introduce the position operator $Q$ and momentum operator $P$ in the rotating frame via
@@ II.2 Renormalized master equation
Figure 1: (a) Quasienergy landscape in phase space of a driven Duffing oscillator. The extrema correspond to the high-amplitude stable state (M) and the low-amplitude stable state (m), while the saddle point (s) marks an unstable state. (b) Cross-section of the quasienergy potential at $P=0$. Quantum energy levels close to the maximum (M) and minimum (m) of the potential are depicted by blue and red lines, respectively. The unstable saddle point (s) is also indicated.
The characteristic behavior of the driven Duffing system is the bistability manifesting as two stable states: the low-amplitude state (LAS) and the high-amplitude state (HAS). As depicted in Fig. 1(a), these stable states correspond to the extrema in the quasienergy landscape, which are defined as the attractors in the phase space. The unstable state, known as the saddle point, is located on the separatrix, serving as the boundary dividing the basins of the attractors. When the damping is present, the system evolves towards the nearby attractor if its initial state lies within the basin of the attractor. However, due to thermal noise, the system does not remain exactly on the attractor, but forms a probability distribution within its basin of attraction. To describe the dissipative dynamics of our system, we employ the Lindblad form master equation
@@
To study the quantum dynamics near the bottom of each stable state,vwe first transform the system to the center of the attractor using the displacement operator
@@
with ${\tilde{\alpha}}=[\frac{1}{2}\kappa+i(\delta\omega+\chi+2\chi|\alpha|^{2})]\alpha+i\varepsilon.$ By choosing $\alpha$ such that $\tilde{\alpha}=0$, the master equation is simplified into
@@
which has the transformation property $S^{\dagger}aS=va+ua^{\dagger}$ with $v=\cosh(|\xi|)$ and $u=-\frac{|\xi|}{\xi}\sinh(|\xi|)$. By defining the squeezed density operator $\bar{\rho}=S^{\dagger}\tilde{\rho}S$, we transform the master equation Eq. (10) into the following form
@@ II.3 Orders of perturbative parameters
To perform perturbation calculations for the renormalized Hamiltonian (II.2), we can choose the dimensionless Planck constant $\lambda=-\chi/(4\Delta)$ as the natural choice for the perturbation parameter. However, since the displacement parameter $\alpha$ is also a function of $\lambda$, it is subtle to properly organize the perturbative terms according to their respective orders. In fact, the Hamiltonian (II.2) should be written in different forms for different attractors. The stable state of the driven Duffing oscillator can be approximated as a coherent state $|\alpha\rangle$. By applying the variational principle in quantum mechanics, i.e., $\partial_{\alpha}\langle\alpha|H|\alpha\rangle=0$, we obtain two solutions for the steady coherent number: a smaller one $|\alpha_{l}|^{2}\approx\beta/(2\lambda)$ for the LAS and a larger one $|\alpha_{h}|^{2}\approx 1/(2\lambda)$ for the HAS.
@@
where the four terms on the right-hand side are given by
@@
$\frac{\bar{H}^{h}}{\delta\omega}$ $=$ $h^{h}_{0}+\gamma h^{h}_{1}+\gamma h^{h}_{2},$ (26)
where the three terms on the right-hand side are
@@
The behavior near the bottom of the LAS is relatively simple and can be modeled using a harmonic oscillator. However, for the HAS, the nonlinear term $\chi(a^{\dagger}a)^{2}\approx\chi|\alpha_{h}|^{2}$ becomes prominent, and the oscillator behaves as a highly squeezed coherent state. In the following sections, we will apply our perturbative method to calculate the crucial quantities related to the HAS of the driven Duffing oscillator, namely, the level spacings, the orbital displacement, and the effective temperature in the vicinity of the HAS attractor.
@@ III.1 Quantum dynamics of HAS
The quantum dynamics of the driven Duffing oscillator near the HAS attractor exhibit a rich interplay among nonlinearity, quantum fluctuations, and thermal noise. In this section, we discuss the quantum properties of the HAS using the renormalized master equation combined with a refined perturbation theory.
@@ III.1.1 Level Spacing
Figure 2: Comparison of perturbation calculations and exact results for energy level spacing. The exact results (lowest curve) for level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ are compared with the zeroth-order (constant spacing), the second-order (second highest line, $\gamma^{2}=\lambda$) and the fourth-order (third highest line, $\gamma^{4}=\lambda^{2}$) corrections. Excellent agreement is observed for low-level numbers near the potential well bottom, while higher-order corrections are required for large $n$. Parameters: $\lambda=0.016$ and $\beta=4/75$.
@@
The nonlinear term $\chi(a^{\dagger}a)^{2}$ in the Hamiltonian, having the opposite sign to $\delta\omega$, results in a decrease in the level spacing $\Delta E_{n}=|E_{n+1}-E_{n}|$ as we approach the saddle point, as illustrated in Fig. 1(b). One can calculate the level spacings with standard perturbation theory by treating the sum of $\gamma h_{1}^{h}$ and $\gamma h_{2}^{h}$ in Eq. (26) as one perturbation term. However, it becomes a challenge to control the accuracy of the level spacings using the perturbative parameter. We find it necessary to distinguish between these two perturbative terms in the perturbation calculations to accurately determine the level spacing. To address this, we have developed a double perturbation theory framework that is particularly suited for the HAS Hamiltonian containing second-order small terms, see the details in Appendix A.
@@
In Fig. 2, we compare our perturbation calculations with the exact numerical results obtained by diagonalizing the original Hamiltonian Eq. (2). These results show an excellent agreement for energy levels near the bottom of the potential well. Under the zeroth-order perturbation approximation, the energy level spacing remains constant across all levels, similar to that of the harmonic oscillator. The second-order and fourth-order corrections provide accuracies up to $\gamma^{2}=\lambda$ and $\gamma^{4}=\lambda^{2}$, respectively. Higher-order perturbation calculations become necessary for levels farther from the bottom.
@@
Figure 3: Orbitial displacement. The average position $\langle N|a|N\rangle$ of the energy level $N$, illustrates the shift due to the quantum fluctuation. Under the harmonic approximation, $\langle N|a|N\rangle$ is a constant for every level (lowest line). Considering higher-order corrections, we observe changes in $\langle N|a|N\rangle$ across levels. Compared with the second-order perturbation result (highest line), the third-order result (second highest line) already aligns well with the exact one (bold line).
III.1.2 Orbital displacement
@@
We denote the eigenstate of the renormalized Hamiltonian $\bar{H}$ as $|n^{\prime}\rangle$, which is generally a superposition of harmonic oscillator eigenstates $|n^{\prime}\rangle=|n\rangle+\sum_{k\neq n}\xi_{kn}|k\rangle$, where the superposition coefficients $\xi_{kn}$ are provided in the Appendix A. The eigenstate $|N\rangle$ of the original Hamiltonian (2) is related to that of the renormalized Hamiltonian (II.2) via the relationship $|N\rangle=DS|n^{\prime}\rangle$. The matrix element $\langle N|a|M\rangle$ for different levels $|N\rangle$ and $|M\rangle$ is then given by
@@
The matrix element $\langle N|a|N\rangle$ provides insight into the orbital displacement in the phase space. Under the harmonic approximation ($\xi_{kn}=0$), $\langle N|a|N\rangle$ remains a constant $\alpha$ for all levels. However, considering higher-order corrections, the orbital displacement $\langle N|a|N\rangle$ changes with energy level, as depicted in Fig. 3. The perturbation results agree well with numerical calculations.
Figure 4: Comparison of stationary probability distributions. (a) Comparison of the stationary probability distribution obtained through perturbation theory to zeroth-order (red line), second-order (blue lines) and exact numerical simulations (black line). (b) Relative error $\Delta p_{n}/p_{n}$ in the stationary distribution obtained from perturbation theory compared to numerical results. The discrepancy increases for higher levels but remains less than 1.
@@
Figure 6: Effective temperature of energy levels. The stationary probability distributions $N_{eff}(n)$ of level $|n\rangle$ obtained through perturbation theory to zeroth order (red line) and second order (blue lines) are compared with the exact numerical simulations (black line). The zero-order effective temperature term gives a constant effective temperature. The second-order term results in changes in $N_{eff}(n)$, which are quite accurate for levels near the bottom.